OpenStudy (anonymous):

A bowler believes he bowls a better score during his second game than his first game 74% of the time. Assuming he is right, what is the probability that in his next 25 pairs of games, he will be correct 18 to 20 times, inclusive? a) Determine the exact probability using the binomial distribution. b) Determine the probability using the normal approximation and compare your answer to the answer from part a).

OpenStudy (anonymous):

\[ \Pr(18\leq X\leq 20) \]

OpenStudy (anonymous):

It seems like a binomial distribution, with 25 trials.

OpenStudy (anonymous):

\[ \Pr(X=k) = \binom{25}{k}(0.74)^k(0.16)^{25-k} \]

OpenStudy (anonymous):

This comes from: \[ \Pr(X=k) = \binom{n}{k}p^k(1-p)^{n-k} \]Where \(n=25\) and \(p=0.74\).

OpenStudy (anonymous):

Since it's a discreet probability distribution:\[ \Large \Pr(18 \leq X \leq 20) = \sum_{k=18}^{20}\Pr(X=k) \]

OpenStudy (anonymous):

@hager Get it?

OpenStudy (kropot72):

@wio ....discrete probability.........

OpenStudy (anonymous):

???

OpenStudy (kropot72):

@hager wio is correct regarding section a) b)The binomial distribution can be approximated by the normal distribution in this case where: \[mean=n \times p\] and \[standard\ deviation=\sqrt{np(1-p)}\]

OpenStudy (anonymous):

thnx guys

OpenStudy (kropot72):

@wio Note that in your first equation (1 - p) should be 0.26 and not 0.16.

OpenStudy (anonymous):

ok thnx

OpenStudy (anonymous):

wio i didn't get for the K

OpenStudy (kropot72):

@hager k is the number of times the player's assumption (that the second game will be better than the first game) is correct. In this problem k=18 or 19 or 20 in the series of calculations.

OpenStudy (anonymous):

so i have to calculate them by these three numbers

OpenStudy (kropot72):

Yes. The first calculation is: \[P(correct\ for\ 18\ games)=\left(\begin{matrix}25 \\ 18\end{matrix}\right)0.74^{18}0.26^{7}=you\ can\ calculate\]

OpenStudy (anonymous):

ok but i have only one result here not three diffrent answers a) 50.33% and b) 49.58%

OpenStudy (anonymous):

but i don't know how

OpenStudy (kropot72):

You will get the correct answer to a) by adding together the 3 values of probability obtained for k = 18, k = 19 and k = 20.

OpenStudy (anonymous):

ok i got it thnx so much

OpenStudy (kropot72):

You're welcome :)